CHIRPED-PULSE  FOURIER  TRANSFORM  MICROWAVE  SPECTROSCOPY  OF  3-

VINYLBENZALDEHYDE

\underline{MIRANDA SMITH} and GORDON G. BROWN, Department of Science and Mathematics, Coker College, 300 E College Ave., Hartsville, SC 29550..

The pure rotational spectrum of 3-vinylbenzaldehyde (3VBA) has been measured and assigned. Coker College’s chirpedpulse Fourier transform microwave (CP-FTMW) spectrometer was used to measure the rotational spectrum of 3VBA in the 7.5 - 18.5 GHz region of the microwave spectrum. The results have been analyzed to discover the rotational constants and centrifugal distortion constants of four distinct conformations of 3VBA: cis,cis-, cis,trans-, trans,cis-, and trans,trans-3VBA. The experimental rotational constants have been compared to the results of ab initio calculations. The performance of Coker’s CP-FTMW spectrometer will also be discussed.